OpenStudy (anonymous):

A hemisphere of radius 3 sits on a horizontal plane. A cylinder stands with its axis vertical, the center of its base at the center of the sphere, and its top circular rim touching the hemisphere. Find the radius and height of the cylinder of maximum volume.

OpenStudy (amistre64):

if heard langrange multipliers are good for this

OpenStudy (amistre64):

 or it has something to do with this kind of picture

OpenStudy (amistre64):

v = pi r^2 h r = 3cos(t) ; h = 3sin(t) v = pi (3cos(t))^2 3sin(t) might be useful

OpenStudy (amistre64):

v' = 27pi(-2cos(t)sin^2(t) + cos^3(t)) if i see it right

OpenStudy (amistre64):

cos(t)(-2sin^2(t)+cos^2(t)) = 0 when t = pi/2, -pi/t which are impractical -2sin^2(t)+cos^2(t) = 0 will be the only things that might make sense

OpenStudy (amistre64):

-2s^2+c^2 = 0 c^2 = 2s^2 1/2 = tan^2(t) 1/sqrt(2) = tan(t); t = arctan(1/sqrt(2))

OpenStudy (amistre64):